# Fast and Tunable Decoherence-Free Subspace Engineering

### v1

@@ V.3.1 Cavity Drive Modification
We start by providing the shortcut Hamiltonian that is used for the preparation of $\lvert\psi_{N,0,0}\rangle$ and $\lvert\psi_{0,0,N}\rangle$, which are both in one-dimensional DFSs. This Hamiltonian can be created by changing the driving amplitude of the injected cavity drive $\eta$ to $\eta+\eta_{s}$ so that $\chi$ transforms to $\chi+\chi_{s}$. Remarkably, this does not require adding any new lasers to the system. We can write this as effectively shifting the Hamiltonian $\hat{H}_{\mathrm{at}}\to\hat{H}_{\mathrm{at}}+\hat{H}_{s}$ where
@@
where the $+$ case will shortcut $\vec{k}=(0,0,N)$ and the -- case applies to $\vec{k}=(N,0,0)$. We assume that $\mu(t)=\beta t$ so $\partial_{t}\mu=\beta$ is constant. In this case, the speed of the shortcut just depends on the maximum $\beta$. As $\mu$ approaches 1, $\chi_{s}$ diverges, indicating that we cannot shortcut all the way to $\mu=1$, and we are limited by laser power. To avoid this divergence, we pick some maximum cutoff value for $|\chi+\chi_{s}|$ as further discussed in Appendix E.1.
@@ V.4 Results
In this section we discuss all the numerical results obtained by simulating the protocols described in the previous sections. To be specific, we simulated the time evolution of the system using a Runge–Kutta algorithm in Python. The dimension of matrices representing operators and states is given by Eq. (14). We note that this detuning may be set to zero without fundamentally changing the system dynamics. We compare our three methods of time evolution for a given DFS. Figure 4 shows results for the single-dimensional DFS $\mathcal{H}_{\text{DFS}}^{(-N)}(\mu)$. Figure 5 uses the three-dimensional DFS $\mathcal{H}_{\text{DFS}}^{(0)}(\mu)$. In both figures, the results of the shortcut method are in solid purple lines, the quench is in dashed blue, and the ramp is graphed in dotted orange.
V.4.1 Edge DFS
First, we consider the extremal value of $C=-N$, i.e. the edge DFS $\mathcal{H}_{\text{DFS}}^{(-N)}(\mu)$. This time-evolution is shown in Fig. 4. In Tab. 1 we show the numerical results for the evolution time $t_{f}$, the final purity $\mathcal{P}(t_{f})$, and the final overlap with the $\mu=1$ target subspace $\mathcal{F}(1,t_{f})$. For the quench, we also provide the final overlap with the $\mu(t_{f})$ DFS $\mathcal{F}(\mu,t_{f})$ because $\mu(t_{f})=\mu_{q}=0.96$ in this case. For the other methods, $\mu(t_{f})=1$, so this final overlap is identical to $\mathcal{F}(1,t_{f})$.
Method $t_{f}(1/\Gamma_{c})$ $\mathcal{P}(t_{f})$ $\mathcal{F}(1,t_{f})$ $\mathcal{F}(\mu(t_{f}),t_{f})$
@@ V.4.1 Edge DFS
Figure 5: Time evolution of $\mathcal{H}_{\text{DFS}}^{(0)}(\mu)$ in units of $t_{f}$ via shortcut (purple, solid), quench (blue, dashed), and ramp (orange, dotted) methods. We shortcut to $\lvert\psi_{N,0,0}\rangle$. Here, $N=5$ and $\Delta_{c}^{\prime}/\kappa=0.1$. The values of $t_{f}$ differ for each method and are found in Tab. 2. (a) Purity of $\hat{\rho}_{\mathrm{at}}(t)$. The quench causes an immediate drop in purity similarly to Fig. 4. The inset shows behavior near $t=t_{f}$. (b) Overlap of $\hat{\rho}_{\mathrm{at}}(t)$ with $\mathcal{H}_{\text{DFS}}^{(0)}(1)$ which measures how well the state matches the target DFS. (c) Probability distribution within $\mathcal{H}_{\text{DFS}}^{(0)}(1)$ for final state $\hat{\rho}_{\mathrm{at}}(t_{f})$. The values of $t_{f}$ are different for the different protocols and given in Tab. 2.
V.4.2 Central DFS
Now, we consider $C=0$, i.e. the central DFS $\mathcal{H}_{\text{DFS}}^{(-N)}(\mu)$. This time-evolution is shown in Fig. 5. Now, we consider time dynamics where we set $C=0$, corresponding to $\mathcal{H}_{\text{DFS}}^{(0)}(\mu)$, as shown in Fig. 5. In Tab. 2 we provide a similar set of numerical results for the previous state in Tab. 1. Here, the quench value for $\mu$ is $\mu(t_{f})=\mu_{q}=0.98$.
Method $t_{f}(1/\Gamma_{c})$ $\mathcal{P}(t_{f})$ $\mathcal{F}(1,t_{f})$ $\mathcal{F}(\mu(t_{f}),t_{f})$
@@ V.4.3 Final State
VI Conclusion and Outlook
@@
We have analyzed an experimentally relevant atom-cavity system wherein one may create and control a DFS. The DFS is in general multi-dimensional, containing degenerate states which do not decohere under photon loss. This DFS comes from diagonalizing the non-Hermitian jump operator via a non-unitary transformation. This diagonalization yields degenerate eigenspaces of the jump operator, where we can tune the external drive to make any of these eigenspaces into a DFS. Within this DFS we show several protocols to prepare pure states by quenching, adiabatically changing the DFS in time, or by employing adiabatic shortcuts. These adiabatic shortcuts allow states to be prepared on rapid timescales only limited by the lifetime of photons in the cavity–i.e. the validity of the model.
@@
Our work provides an important stepping stone towards controlling multi-dimensional DFSs in an atom-cavity system, and opens the door to several interesting questions. Future work will explore the use of this system in experimentally relevant parameter regimes, where we expect to see protection against experimentally relevant noise sources. Of particular interest is the control of the dimensionality of the DFS in order to create qudits which are protected against photon decay. These protected qudits may be relevant for quantum computing and quantum metrology [66, 67]. The primary challenge would be defining logical operations which do not couple states out of the DFS, but are experimentally feasible.
@@
VII Acknowledgments
We would like to thank Shah S. Alam and Nicholas T. Weir for useful discussions. This work is supported by the Department of Energy Quantum Systems Accelerator, Grant No. 7565477, and the National Science Foundation (Grants No. PHY-2317149 and No. OSI-2016244).
@@ A.1 Interaction Picture
$\hat{H}^{\prime}=$ $\hbar\frac{\omega_{1}+\omega_{2}}{2}\hat{a}^{\dagger}\hat{a}+\sum_{j=1}^{N}\frac{-\omega_{1}+\omega_{2}}{2}\hbar\left(\lvert 1_{g}\rangle_{j}\langle 1_{g}\rvert-\lvert\text{-}1_{g}\rangle_{j}\langle\text{-}1_{g}\rvert\right)$ (46)
@@
A.2 Excited State Elimination
@@
A.3 Cavity Mode Elimination
@@ Appendix E Shortcut Derivation
[24] for any $\vec{n}$ such that $n_{3}-n_{1}\neq C=k_{3}-k_{1}$, meaning all $\lvert\psi_{\vec{n}}^{\perp}\rangle\in\mathcal{H}_{\text{CS}}$. We may apply this additional Hamiltonian in concert with the original $\hat{H}_{\mathrm{at}}$ in order to drive directly from the initial state $\lvert\psi(0)\rangle$ into the target state.
@@
$+\mathcal{N}_{\vec{k}}(\hat{c}_{1}^{\dagger})^{k_{1}}(\hat{c}_{2}^{\dagger})^{k_{2}}[k_{3}(\hat{c}_{3}^{\dagger})^{k_{3}}\partial_{t}\hat{c}_{3}^{\dagger}]\lvert 0\rangle.$
@@
with $n_{p}=k_{p}+1$ corresponding to a vector in Eq. (90). Therefore, a shortcut Hamiltonian must couple $|\psi_{\vec{k}}\rangle$ to exactly the $|\psi_{\vec{n}}\rangle$ where $\vec{n}$ satisfies Eq. (90).
E.1 Cavity Drive Modification
@@
We can apply this to $\lvert\psi_{\vec{k}}\rangle$ and find that $\langle\psi_{\vec{n}}^{\perp}\rvert\hat{H}_{s}\lvert\psi_{\vec{k}}\rangle$ is nonzero for exactly the $\vec{n}$ given in Eq. (90) and define $p\in\{1,2,3\}$ such that $n_{p}=k_{p}+1$. The result for these vectors reads

### v2

@@ V.3.1 Cavity drive modification
We start by providing the shortcut Hamiltonian that is used for the preparation of $\lvert\psi_{N,0,0}\rangle$ and $\lvert\psi_{0,0,N}\rangle$, which are both in one-dimensional DFSs. This Hamiltonian can be created by changing the driving amplitude of the injected cavity drive $\eta$ to $\eta+\eta_{s}$ so that $\chi$ transforms to $\chi+\chi_{s}$. Remarkably, this does not require adding any new lasers to the system. We can write this as effectively shifting the Hamiltonian $\hat{H}_{\mathrm{at}}$ to $\hat{H}_{\mathrm{at}}+\hat{H}_{s}$ where
@@
where the $+$ case will shortcut $\vec{k}=(0,0,N)$ and the -- case applies to $\vec{k}=(N,0,0)$. We assume that $\mu(t)=\beta t$ so $\partial_{t}\mu=\beta$ is constant. In this case, the speed of the shortcut just depends on the maximum $\beta$. As $\mu$ approaches 1, $\chi_{s}$ diverges, indicating that we cannot shortcut all the way to $\mu=1$, and we are limited by laser power. Note that this divergence cannot be remedied by choosing a different $\mu(t)$ dependence. To avoid the issue, we pick some maximum cutoff value for $|\chi+\chi_{s}|$ beyond which we continue ramping $\mu$ but keep $|\chi+\chi_{s}|$ constant. This has a minimal effect on the characteristics of the evolution overall, and is further discussed in Appendix E.1.
@@ V.4 Results
In this section we discuss all the numerical results obtained by simulating the protocols described in the previous sections. To be specific, we simulated the time evolution of the system using a Runge–Kutta algorithm in Python. The dimension of matrices representing operators and states is given by Eq. (14). We compare our three methods of time evolution for a given DFS. Figure 4 shows results for the single-dimensional DFS $\mathcal{H}_{\text{DFS}}^{(-N)}(\mu)$. Figure 5 uses the three-dimensional DFS $\mathcal{H}_{\text{DFS}}^{(0)}(\mu)$. In both figures, the results of the shortcut method are in solid purple lines, the quench is in dashed blue, and the ramp is graphed in dotted orange.
V.4.1 Edge decoherence-free subspace
First, we consider the extremal value of $C=-N$, i.e. the edge DFS $\mathcal{H}_{\text{DFS}}^{(-N)}(\mu)$. This time-evolution is shown in Fig. 4. In Tab. 1 we show the numerical results for the evolution time $t_{f}$, the final purity $\mathcal{P}(t_{f})$, and the final overlap $\mathcal{F}(1,t_{f})$ with the $\mu=1$ target subspace. For the quench, we also provide the final overlap $\mathcal{F}(\mu,t_{f})$ with the $\mu(t_{f})$ DFS because $\mu(t_{f})=\mu_{q}=0.96$ in this case. For the other methods, $\mu(t_{f})=1$, so this final overlap is identical to $\mathcal{F}(1,t_{f})$.
Method $t_{f}/\Gamma_{c}$ $\mathcal{P}(t_{f})$ $\mathcal{F}(1,t_{f})$ $\mathcal{F}(\mu(t_{f}),t_{f})$
@@ V.4.1 Edge decoherence-free subspace
Figure 5: Time evolution of $\mathcal{H}_{\text{DFS}}^{(0)}(\mu)$ in units of $t_{f}$ via shortcut (purple, solid), quench (blue, dashed), and ramp (orange, dotted) methods. We shortcut to $\lvert\psi_{0,N,0}\rangle$. Here, $N=5$ and $\Delta_{c}^{\prime}/\kappa=0.1$. The values of $t_{f}$ differ for each method and are found in Tab. 2. (a) Purity of $\hat{\rho}_{\mathrm{at}}(t)$. The quench causes an immediate drop in purity similarly to Fig. 4. The inset shows behavior near $t=t_{f}$. (b) Overlap of $\hat{\rho}_{\mathrm{at}}(t)$ with $\mathcal{H}_{\text{DFS}}^{(0)}(1)$ which measures how well the state matches the target DFS. (c) Probability distribution within $\mathcal{H}_{\text{DFS}}^{(0)}(1)$ for final state $\hat{\rho}_{\mathrm{at}}(t_{f})$. The values of $t_{f}$ are different for the different protocols and given in Tab. 2.
@@
V.4.2 Central decoherence-free subspace
Now, we consider time dynamics where we set $C=0$, corresponding to $\mathcal{H}_{\text{DFS}}^{(0)}(\mu)$, as shown in Fig. 5. In Tab. 2 we provide a similar set of numerical results for the previous state in Tab. 1. Here, the quench value for $\mu$ is $\mu(t_{f})=\mu_{q}=0.98$.
Method $t_{f}/\Gamma_{c}$ $\mathcal{P}(t_{f})$ $\mathcal{F}(1,t_{f})$ $\mathcal{F}(\mu(t_{f}),t_{f})$
@@ V.4.3 Final state
VI Realizing a Qubit
Once we have prepared a DFS, we may drive between decoherence-free states without causing any decoherence. This allows us to store quantum information and perform quantum information tasks by building a logical qubit which is stabilized by cavity-photon loss. Here, we focus on a two-dimensional DFS at $\mu=1$ for simplicity, but these results may be generalized.
A DFS will satisfy $\mathfrak{D}(N,C)=2$ exactly when
$C\in\{\pm(N-3),\pm(N-2)\}$ (45)
for $N\geq 3$, and $C=0$ for $N=2$. This can be derived using Eq. (29). Denote the two states in this subspace as
$|g\rangle=$ $|\psi_{k_{1},k_{2},k_{3}}\rangle$ (46)
$|e\rangle=$ $|\psi_{k_{1}+1,k_{2}-2,k_{3}+1}\rangle$
for fixed $k_{1},k_{2},k_{3}$ such that $C=k_{3}-k_{1}$ satisfies Eq. (45). As each multi-dimensional DFS in our system contains entangled states, the operators which drive between $|g\rangle$ and $|e\rangle$ must contain multi-particle interaction terms. Any single-particle term such as $\hat{c}_{i}^{\dagger}\hat{c}_{j}$ for $i\neq j$ will necessarily move some population outside of the DFS by changing $k_{3}-k_{1}$. For an operator to drive between states within a fixed DFS without causing decoherence, a sufficient condition is to enforce that the operator commutes with $\hat{L}$ [67]. In this case, any time-evolution caused by such an operator will commute with $\hat{L}$, meaning that if the initial state of the system is a dark state of $\hat{L}$, the final state must be as well.
To drive between qubit states, we define effective Pauli operators [66] as follows:
$\hat{\sigma}_{x}=$ $\frac{1}{\mathcal{N}(N,C)}(\hat{c}_{2}^{\dagger}\hat{c}_{3}\hat{c}_{2}^{\dagger}\hat{c}_{1}+\hat{c}_{1}^{\dagger}\hat{c}_{2}\hat{c}_{3}^{\dagger}\hat{c}_{2})$ (47)
$\hat{\sigma}_{y}=$ $\frac{i}{\mathcal{N}(N,C)}(\hat{c}_{2}^{\dagger}\hat{c}_{3}\hat{c}_{2}^{\dagger}\hat{c}_{1}-\hat{c}_{1}^{\dagger}\hat{c}_{2}\hat{c}_{3}^{\dagger}\hat{c}_{2})$
$\hat{\sigma}_{z}=$ $\frac{-i}{2}\left[\hat{\sigma}_{x},\hat{\sigma}_{y}\right]$
$=$ $\frac{1}{\mathcal{N}^{2}(N,C)}(4\hat{n}_{1}\hat{n}_{2}\hat{n}_{3}-\hat{n}_{1}\hat{n}_{2}^{2}-\hat{n}_{2}^{2}\hat{n}_{3}$
$-\hat{n}_{2}^{2}+2\hat{n}_{1}\hat{n}_{3}+\hat{n}_{1}\hat{n}_{2}+\hat{n}_{2}\hat{n}_{3}+\hat{n}_{2})$
where $\hat{n}_{i}=\hat{c}_{i}^{\dagger}\hat{c}_{i}$ for $i\in\{1,2,3\}$ and
$\mathcal{N}(N,C)=\begin{cases}\sqrt{6N-12}&|C|=N-3,\\ \sqrt{2N-2}&|C|=N-2,\end{cases}$ (48)
is a normalization dependent on the DFS. Note that these operators are Hermitian only at $\mu=1$ where the $\hat{c}_{i}$ operators satisfy canonical commutation relations.
The above operators behave identically to typical Pauli operators when acted on states within the chosen DFS. For complex coefficients $\alpha,\beta$, one can verify
$\hat{\sigma}_{x}\left(\alpha|g\rangle+\beta|e\rangle\right)=$ $\beta|g\rangle+\alpha|e\rangle,$ (49)
$\hat{\sigma}_{y}\left(\alpha|g\rangle+\beta|e\rangle\right)=$ $i\beta|g\rangle-i\alpha|e\rangle,$
$\hat{\sigma}_{z}\left(\alpha|g\rangle+\beta|e\rangle\right)=$ $-\alpha|g\rangle+\beta|e\rangle,$
which follows the usual operation of Pauli matrices.
Note that these operators behave differently from Pauli operators when the entire Hilbert space is considered. In particular, Pauli operators satisfy
$\left[\hat{\sigma}_{i},\hat{\sigma}_{j}\right]=2i\epsilon_{i\,j\,k}\hat{\sigma}_{k}$ (50)
for $i,j,k\in\{x,y,z\}$ where $\epsilon_{i\,j\,k}$ is the Levi-Civita symbol [66]. In this system, this relation holds only when acting on states within the DFS and is not true over the entire Hilbert space because in general, $\left[\hat{\sigma}_{i},\hat{\sigma}_{j}\right]$ contains higher orders of the $\hat{c}_{i}$ operators.
The entangling operations described by the $\hat{\sigma}_{i}$ must be engineered, but can in principle be realized in experiment [14]. Moreover, this complexity comes with a great advantage: the very same complexity that is needed to control the qubit also hinders flips between the states $\lvert g\rangle$ and $\lvert e\rangle$. The reason for this is that the qubit is protected by the DFS, so such flips require two-particle terms which makes the qubit naturally robust against single-particle decoherence.
VII Conclusion and Outlook
@@
We have analyzed an experimentally relevant atom-cavity system wherein one may create and control a DFS. The DFS is in general multi-dimensional, containing degenerate states which do not decohere under photon loss. This DFS comes from diagonalizing the non-Hermitian jump operator via a non-unitary transformation. This diagonalization yields degenerate eigenspaces of the jump operator, where we can tune the external drive to make any of these eigenspaces into a DFS. Within this DFS we show several protocols to prepare pure states by quenching, adiabatically changing the DFS in time, or by employing adiabatic shortcuts. These adiabatic shortcuts allow states to be prepared on rapid timescales only limited by the lifetime of photons in the cavity–i.e. the validity of the model. We also provide operators which fully control a decoherence-free qubit without causing any decoherence.
@@
Our work provides an important stepping stone towards controlling multi-dimensional DFSs in an atom-cavity system, and opens the door to several interesting questions. Future work will explore the use of this system in experimentally relevant parameter regimes, where we expect to see protection against experimentally relevant noise sources. Of particular interest is the control of the dimensionality of the DFS in order to create qudits which are protected against photon decay. These protected qudits may be relevant for quantum computing and quantum metrology [68, 69].
@@
VIII Acknowledgments
We would like to thank Shah Saad Alam and Nicholas T. Weir for useful discussions. This work is supported by the Department of Energy Quantum Systems Accelerator (Grant No. 7565477) and the National Science Foundation (Grants No. PHY-2317149 and No. OSI-2016244).
@@ A.1 Interaction picture
$\hat{H}^{\prime}=$ $\hbar\frac{\omega_{1}+\omega_{2}}{2}\hat{a}^{\dagger}\hat{a}+\sum_{j=1}^{N}\frac{-\omega_{1}+\omega_{2}}{2}\hbar(\lvert 1_{g}\rangle_{j}\langle 1_{g}\rvert-\lvert\text{-}1_{g}\rangle_{j}\langle\text{-}1_{g}\rvert)$ (52)
@@
A.2 Excited-state elimination
@@
A.3 Cavity-mode elimination
@@ Appendix E Shortcut Derivation
[24] for any $\vec{n}$ such that $n_{3}-n_{1}\neq C=k_{3}-k_{1}$, meaning all $\lvert\psi_{\vec{n}}^{\perp}\rangle\in\mathcal{H}_{\text{CS}}^{(C)}(\mu)$. We may apply this additional Hamiltonian in concert with the original $\hat{H}_{\mathrm{at}}$ in order to drive directly from the initial state $\lvert\psi(0)\rangle$ into the target state.
@@
$+\mathcal{N}_{\vec{k}}(\hat{c}_{1}^{\dagger})^{k_{1}}(\hat{c}_{2}^{\dagger})^{k_{2}}[k_{3}(\hat{c}_{3}^{\dagger})^{k_{3}-1}\partial_{t}\hat{c}_{3}^{\dagger}]\lvert 0\rangle.$
@@
where $p\in\{1,2,3\}$ is such that $n_{p}=k_{p}+1$. Therefore, a shortcut Hamiltonian must couple $|\psi_{\vec{k}}\rangle$ to exactly the $|\psi_{\vec{n}}\rangle$ where $\vec{n}$ satisfies Eq. (96).
E.1 Cavity-drive modification
@@
We can apply this to $\lvert\psi_{\vec{k}}\rangle$ and find that $\langle\psi_{\vec{n}}^{\perp}\rvert\hat{H}_{s}\lvert\psi_{\vec{k}}\rangle$ is nonzero for exactly the $\vec{n}$ given in Eq. (96). We again define $p\in\{1,2,3\}$ such that $n_{p}=k_{p}+1$. The result for these vectors reads